# Two-Sample Testing with Block-Wise Missingness in Multi-Source Data

### v1

@@ 4 Simulation Studies
      Location and scale: $t_{5}(\mu\mathbf{1}_{d},\Sigma_{Y})$, with $\mu=0.4\log{d}/\sqrt{{d}}$ and $\Sigma_{Y,ij}=0.8^{|i-j|}$.
We set the default setting as $m=n=100$, $L=2$, $p_{X}=p_{Y}=0.5$, and $d\in\{200,500,1000\}$. To validate the performance of tests under diverse scenarios, we also include additional experiments on power of tests with imbalanced sample sizes ($m=200,n=100$), increased numbers of sources ($L=3$), varied sampling rates ($p_{X}=p_{Y}\in\{0.2,0.8\}$) and imbalanced sampling rates ($p_{X}=0.5,p_{Y}=0.8$) in Supplementary O. We repeat $1000$ times for simulations of empirical sizes and $100$ times for simulations of power of tests.
At significance level $\theta=0.05$, the empirical sizes under both balanced and imbalanced sampling rates under different groups are reported in Table 1. As shown in the table, when the sampling rates are balanced, all methods have acceptable type-I error. But when the sampling rates are imbalanced, BRISE-c, BRISE-v, MMD-Miss and methods using complete cases still control type-I error well, while BRISE-c(sp) fails to control the type I error. This highlights the advantage of our methods on considering block-wise missingness mechanism.
@@
The estimated powers of the tests under the default setting are provided in Table 2. Across all dimensions, the BRISE family consistently achieves superior statistical power compared to the baselines. In particular, BRISE-c demonstrates the most robust performance, maintaining high power across a broad range of distributional differences.
Table 2: Estimated powers at significance level $\theta=0.05$, for $m=n=100$, $L=2$ and $p_{X}=p_{Y}=0.5$. Under the multivariate Gaussian I: a. Location, b. Scale, and c. Mixed; the multivariate log-normal II: a. Location, b. Scale, and c. Mixed; and the multivariate $t_{5}$ III: a. Location, b. Scale, and c. Mixed. Top-1, Top-2 powers are in bold and underlined, respectively.
@@
RISE 55 33 19 39 35 22 42 28 22
MMD 7 1 0 7 12 9 12 1 0
BD 4 6 4 42 41 37 4 6 3
MT 13 11 9 8 11 8 41 24 27
Under the multivariate Gaussian distribution (Setting I), BRISE-c demonstrates consistently superior performance across all three subsettings, achieving the highest power for location, scale, and mixed alternatives. BRISE-v follows closely, showing strong power in most subsettings. In contrast, RISE and BD exhibit moderate power in scale settings but perform weakly in location and mixed cases. MMD shows limited power, with some effectiveness for location alternatives at lower dimensions ($d=200$). MT also exhibits limited power across all subsettings, with slightly better performance for mixed alternatives at higher dimensions ($d=1000$). MMD-Miss fails to show any power across all subsettings, as it is not designed for data with block-wise missingness in high dimensions.
For the multivariate log-normal distribution (Setting II), BRISE-v performs strongly in detecting location differences across all dimensions. BRISE-c dominates in scale and mixed subsettings, achieving the highest power. BD is competitive in scale and mixed settings, especially at higher dimensions, but is outperformed by BRISE-c. RISE, MMD, and MT exhibit moderate power, with RISE and MT performing well for mixed alternatives across all dimensions. MMD-Miss shows no power in this block-wise missingness scenario.
With the heavy-tailed multivariate $t_{5}$ distribution (Setting III), BRISE-v achieves the highest power for location shifts. BRISE-c excels in scale and mixed alternatives, outperforming other methods. RISE shows moderate power across all alternatives. BD exhibits moderate power in the scale setting but low power otherwise. MT shows moderate power in the mixed setting but low power in other alternatives.
Under Setting I, we further study the trend of statistical power of all the methods when $p_{X}$ and $p_{Y}$ vary simultaneously ($p_{X}=p_{Y}\in\{0,0.1,...,0.9,1\}$); see Figure 5. BRISE-c consistently achieves high power across most dimensions and probability values, while BRISE-v also exhibits reasonably high powers; both BRISE variants are robust to variations in sampling rates, with only minor power reduction as sampling rates decrease. In contrast, other methods (MMD-Miss, RISE, MMD, BD, MT) suffer significant power loss under decreasing sampling rates or fail to perform effectively at nearly all sampling rates. Specifically, when sampling rates are high, we observe that MMD performs best under Setting I-a; BD is very powerful under Setting I-b, the scale alternative; while RISE precedes under Setting I-c, the mixed alternative. However, these methods suffer from reduced sample size at low sampling rates, underscoring the robustness of BRISE. MT and MMD-Miss show limited effectiveness across most alternatives in Setting I. Critically, BRISE’s resilience to missingness is consistently observed across other experimental settings beyond Setting I (see Supplementary O), underscoring its general reliability.
@@
In addition, we conduct simulations under other hyperparameter scenarios, with the corresponding results presented in Supplementary O. Specifically, for scenarios with imbalanced sample sizes, they show similar phenomena to the default setting; for scenarios with different sampling probabilities, including both lower and higher cases, we observe that BRISE-family methods are more powerful at lower sampling rates compared to other methods; we also examine scenarios where $p_{X}\neq p_{Y}$, and the simulation power still exhibits similar behaviors as in the default case. Furthermore, we explore settings with an increased number of sources, specifically $L=3$. In this scenario, the power of all BRISE methods remains at the same level or degrades only slightly compared to the two-source case, while other methods using complete-case analysis suffer significant power loss due to smaller sample sizes.
In summary, the simulation results provide strong empirical support for the effectiveness of the BRISE family of tests. Each variant demonstrates a unique strength for certain alternatives. Collectively, utilizing the BPET framework, BRISE significantly outperforms traditional tests, making it well-suited for modern high-dimensional two-sample testing problems.
@@ 5.1 Sepsis Identification Using Multi-source Biomarkers
We now illustrate the utility of the BRISE using a sepsis study. Sepsis is a major public health concern in the United States, characterized by the body’s pathologic response to infection, which can lead to tissue damage, organ failure, and death (Singer et al., 2016). Identifying biomarkers that can provide early warning signs and enable timely, targeted interventions is therefore crucial for the treatment of sepsis. The goal of this study is to examine potential differences in age, sex, race, APACHE II, absolute neutrophil count (ANC), absolute lymphocyte count (ALC), monocyte distribution width (MDW), and other biomarkers among septic patients and patients who were considered critically ill without sepsis (CINS).
Most single studies involving these biomarkers have relatively small sample sizes or focus on patients who present to the emergency department with community-acquired sepsis (Agnello et al., 2021). In contrast, this study evaluates whether biomarker distributions differ between septic patients and CINS in the context of purely hospital-acquired sepsis as observed in surgical and trauma ICUs. This task is particularly challenging and important: patients admitted to surgical/trauma ICUs often experience concomitant inflammation due to surgical or traumatic injury (clinically difficult to distinguish from sepsis) and are at heightened risk of sepsis due to invasive procedures, indwelling devices (e.g., catheters, ventilators), and compromised immune state from prolonged hospitalization or underlying health conditions (Mas-Celis et al., 2021; Cohen et al., 2023).
To investigate this question, we combined data from three prospective observational studies of hospitalized patients admitted to a surgical ICU at a quaternary-care academic health science center. Two studies were conducted by the UF Sepsis and Critical Illness Research Center (SCIRC) and are registered on clinicaltrials.gov (NCT04414189, NCT05110937) (Brakenridge et al., 2021). The third study involved subjects enrolled at Shands-UF Hospital as part of an NIGMS-funded consortium (Barrios et al., 2024). A key challenge in this analysis is the presence of block-wise missingness arising from differences in study protocols. Specifically, in two studies, biomarkers were measured on Days $1$ and $4$, while in the third study they were measured on Day $2$. No single study collected biomarkers on all days, leading to a block-missing pattern in the combined dataset.
@@
Figure 6: Heatmap visualization of sepsis biomarker data. Gray indicates missing values.
Table 3 reports the $p$-values obtained from our method and benchmark approaches. Both BRISE variants detect highly significant differences between septic and non-septic patients. In contrast, MMD-Miss fails to reject the null hypothesis, reflecting its conservative behavior, as it was not designed for this type of setting. Complete-case (CC) analysis is not applicable here, as no complete cases are available. These results underscore the superior performance and robustness of BRISE in real-world biomedical studies with block-wise missingness.
Table 3: The $p$-values of the tests for the sepsis data. Significant values are shown in bold.
@@ 5.2 Multi-modal Differentiation of Alzheimer’s Disease
A critical real-world application of multi-modal statistical analysis is in distinguishing between cognitive health states based on medical data. In particular, Alzheimer’s Disease (AD) and cognitively normal (CN) conditions often exhibit subtle but detectable differences across biomarkers, imaging features, and metabolic indicators. Accurate identification of these differences is crucial for early diagnosis, targeted intervention, and a better understanding of disease mechanisms.
@@
To evaluate the practical effectiveness of our methods, we apply them to the Alzheimer’s Disease Neuroimaging Initiative (ADNI) dataset (Mueller et al., 2005a). Each ADNI participant is uniquely identified by a Participant Roster ID (RID). Following the approach of Balsis et al. (2015), we define diagnostic groups using Mini-Mental State Examination (MMSE) scores (Folstein et al., 1975): individuals with MMSE scores greater than $24$ and up to $30$ are classified as CN, while those with scores of $24$ or below are classified as AD. For individuals with multiple MMSE measurements, we use the average score for consistency.
@@
Our analysis integrates three modalities: Magnetic Resonance Imaging (MRI), Positron Emission Tomography (PET), and biospecimen data. MRI and PET image data are processed into numeric variables by the Center for Imaging of Neurodegenerative Diseases at the University of California, San Francisco, and the Jagust Lab at the University of California, Berkeley, respectively. The biospecimen data, analyzed by the Proteomics and Metabolomics Shared Resource at Duke University, provide serum measurements of selected metabolites, including biomarkers such as carnitine and decanoylcarnitine. All data are available through the Analysis Ready Cohort (ARC) Builder [^1] https://ida.loni.usc.edu/explore/jsp/search_v2/search.jsp?project=ADNI.
@@
We merge the three modalities using RID after removing cases with random missingness. The proposed methods are then evaluated on two datasets: (i) a balanced subset (Sub) consisting of $200$ randomly selected individuals from each group (AD and CN), and (ii) the full dataset (Full) comprising all available samples, including $2,047$ CN and $507$ AD subjects. Each method is assessed using both permutation-based (P) inference ($1000$ permutations) and asymptotic (A) null approximations when available.
@@
Table 4 reports the resulting $p$-values. Both BRISE variants achieve highly significant results across all datasets and inference types, demonstrating superior statistical power even under substantial block-wise missingness. By contrast, RISE, MMD, and BD yield significant results only on the full dataset and fail to reject the null hypothesis in the balanced subset. MMD-Miss produces $p$-values equal to $1$ in all settings, likely due to its inability to leverage cross-pattern structure. MT fails in both datasets, further illustrating its limited robustness for incomplete multi-modal data. These findings align with our simulation studies and highlight the practical applicability of BRISE for detecting group differences in high-dimensional, block-wise incomplete settings.
Table 4: The $p$-values of the tests for the ADNI data. Significant values are shown in bold.
@@ 6 Discussion and Conclusion
In this work, we propose a general framework, called BPET, for two-sample testing under block-wise missingness without requiring imputation or deletion of observations, enabling reliable inference with incomplete datasets. Further, we incorporate the rank-based RISE test within the proposed BPET framework to account for possible multiple modalities. We established the finite-sample properties, validity, and consistency of the proposed methods. Their effectiveness was demonstrated through both extensive simulations and two real-world applications, and the BRISE implementation exhibited strong empirical performance. Although developed for two-sample testing, BPET can be readily extended to other inferential tasks such as independence testing, clustering, and change-point detection in the presence of block-wise missingness.
As a general framework, we emphasize that BPET is not limited to graph/rank-based methods; a broad class of two-sample test procedures can be applied within the BPET framework. For example, by computing pattern-wise dissimilarities or kernel metrics (Section 2.1), both kernel-based, and distance-based methods can be incorporated into BPET. Beyond the $k$-NNG, other similarity graphs such as $k$-MST (Friedman and Rafsky, 1979) and $k$-MDP (Rosenbaum, 2005) may also be used. These approaches follow the same principle: partition the data into missingness patterns, perform pattern-specific computations, and aggregate the results into a global test statistic. The finite-sample properties established in Section 3.1 can be extended, provided the method produces a symmetric similarity matrix such as $\mathbf{R}$.
In multi-modal data, it is common that samples from different patterns have non-overlapping variables (or modalities). In the main text, we proposed the zero-filling strategy in Section 2.1; one may also adopt alternative measures, such as rank-based distances (Supplementary L) or embedding all sources into a latent subspace (e.g., via partial multi-view clustering (Li et al., 2014; Zhao et al., 2016; Wang et al., 2018; Feng et al., 2024)) before computing distances on the aligned projections. An interesting observation from our simulation is that the proposed method leveraging the zero-filling strategy consistently outperforms other dissimilarity measures. This raises a future research question: whether information from samples without shared or overlapping sources/modalities is important, and if so, how to efficiently leverage it. Our work is among the first to explore these important questions.

### v2

@@ 5 Simulation Studies
      Mixed: $t_{5}(\mu\mathbf{1}_{d},\Sigma_{Y})$, with $\mu=0.4\log{d}/\sqrt{{d}}$ and $\Sigma_{Y,ij}=0.8^{|i-j|}$.
We set the default setting as $m=n=100$, $L=2$, $p_{X}=p_{Y}=0.5$, and $d\in\{200,500,1000\}$. We also include additional experiments in Supplementary P, including imbalanced sample sizes ($m=200,n=100$), an increased number of sources ($L=3$), varied sampling rates ($p_{X}=p_{Y}\in\{0.2,0.8\}$), and imbalanced sampling rates ($p_{X}=0.5,p_{Y}=0.8$). We repeat simulations $1000$ times for empirical sizes and power.
At significance level $\theta=0.05$, the empirical sizes under both balanced and imbalanced sampling rates are reported in Table 1. When the sampling rates are balanced, all methods exhibit acceptable type-I error. When the sampling rates are imbalanced, BRISE-c, BRISE-v, MMD-Miss, and the complete-case methods still control type-I error well, whereas BRISE-c(sp) fails to do so. This highlights the importance of accounting for block-wise missingness via pattern-wise permutation.
In the low-dimensional experiments, the imputation-based methods using MICE fail to control the type-I error, despite being computationally feasible in these settings. This suggests that imputation may not provide valid inference under block-wise missingness, even when dimensionality is low.
@@
$d$ 10 20 50 10 20 50 10 20 50
$p_{X}=p_{Y}=0.5$ Setting I Setting II Setting III
BRISE-c 0.062 0.048 0.057 0.045 0.043 0.056 0.040 0.052 0.057
BRISE-v 0.062 0.061 0.049 0.068 0.064 0.055 0.067 0.055 0.048
MMD-Miss 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
RISE 0.040 0.055 0.056 0.044 0.049 0.048 0.048 0.044 0.055
MMD 0.035 0.039 0.017 0.031 0.037 0.016 0.039 0.023 0.024
BD 0.041 0.058 0.043 0.044 0.046 0.056 0.047 0.044 0.057
MT 0.040 0.056 0.042 0.040 0.057 0.052 0.056 0.057 0.037
MICE+RISE 0.906 0.996 1.000 0.867 0.981 1.000 0.914 0.991 1.000
MICE+MMD 0.530 0.729 0.915 0.490 0.627 0.998 0.596 0.746 0.967
MICE+BD 0.311 0.381 0.616 0.217 0.182 0.392 0.303 0.233 0.422
MICE+MT 0.013 0.177 0.854 0.037 0.084 0.433 0.014 0.196 0.886
Table 2: Estimated powers at significance level $\theta=0.05$ with $m=n=100$, $L=2$, and $p_{X}=p_{Y}=0.5$. Results are shown for multivariate Gaussian (I), log-normal (II), and $t_{5}$ (III) distributions under location (a), scale (b), and mixed (c) alternatives. Top-1 and Top-2 powers are shown in bold and underlined, respectively.
@@
RISE 0.446 0.306 0.197 0.326 0.287 0.244 0.416 0.303 0.215
MMD 0.068 0.020 0.009 0.077 0.088 0.077 0.093 0.019 0.007
BD 0.054 0.041 0.041 0.352 0.378 0.386 0.058 0.042 0.043
MT 0.081 0.086 0.069 0.073 0.087 0.074 0.359 0.331 0.273
The estimated powers under the default setting are reported in Table 2. Across all dimensions and alternatives, the BRISE family consistently achieves higher power than the competing methods. In particular, BRISE-c shows the most stable performance, maintaining strong power across a wide range of distributional differences, while BRISE-v remains competitive, especially under heterogeneous alternatives.
Figure 4 further examines power as the sampling rates vary with $p_{X}=p_{Y}\in{0,0.1,\ldots,0.9,1}$ under Setting I. Both BRISE-c and BRISE-v are robust to decreasing sampling rates, with only modest power loss. In contrast, MMD-Miss, RISE, MMD, BD, and MT experience substantial power degradation or fail to perform reliably as sampling rates decrease.
When sampling rates are high, some baselines perform well for specific alternatives (MMD for location, BD for scale, and RISE for mixed alternatives). However, these methods rely on complete-case analysis and suffer from severe effective sample size reduction at low sampling rates, underscoring the robustness advantage of BRISE. MT and MMD-Miss show limited power across most settings.
Additional results under alternative distributions, imbalanced sample sizes, unequal sampling probabilities, and increased numbers of sources are provided in Supplementary P. Across all scenarios, BRISE consistently maintains valid type-I error control and strong power, whereas complete-case methods often exhibit inflated or overly conservative type-I error or substantial power loss, provided Condition (2) holds.
Overall, the simulation studies provide strong empirical support for the BPET framework and demonstrate that the BRISE tests are well suited for high-dimensional two-sample testing with block-wise missing data.
@@ 6.1 Sepsis Identification Using Multi-source Biomarkers
We illustrate the utility of BRISE using a sepsis study. Sepsis is a major public health concern associated with high morbidity and mortality (Singer et al., 2016). The goal of this analysis is to assess whether the distributions of age, sex, race, APACHE II score, absolute neutrophil count (ANC), absolute lymphocyte count (ALC), monocyte distribution width (MDW), and other biomarkers differ between septic patients and critically ill non-septic (CINS) patients.
Most existing studies of these biomarkers involve relatively small samples or focus on community-acquired sepsis in emergency department settings (Agnello et al., 2021). In contrast, we consider hospital-acquired sepsis in surgical and trauma ICUs, where distinguishing sepsis from inflammation due to surgery or trauma is clinically challenging, and where patients are at elevated risk due to invasive procedures, indwelling devices, and prolonged hospitalization (Mas-Celis et al., 2021; Cohen et al., 2023).
The data were combined from three prospective observational studies of patients admitted to a surgical ICU at a quaternary-care academic medical center. Two studies were conducted by the UF Sepsis and Critical Illness Research Center (SCIRC) and registered at clinicaltrials.gov (NCT04414189, NCT05110937) (Brakenridge et al., 2021), and the third study was part of an NIGMS-funded consortium at Shands-UF Hospital (Barrios et al., 2024). A key methodological challenge is block-wise missingness induced by study-specific protocols: biomarkers were measured on Days 1 and 4 in two studies, but on Day 2 in the third, and no study collected all biomarkers across all days. As a result, the combined dataset exhibits block-wise missing patterns that preclude standard complete-case or covariate-adjusted analyses.
@@
Figure 5: Heatmap of sepsis biomarker data. Gray indicates missing values.
Table 3 reports the $p$-values from BRISE and benchmark methods. Both BRISE-c and BRISE-v detect highly significant differences between septic and non-septic patients. In contrast, MMD-Miss fails to reject the null hypothesis, reflecting its conservative behavior in this setting, and complete-case analysis is infeasible due to the absence of fully observed samples. These results demonstrate the practical advantage of BRISE for real-world biomedical studies with block-wise missingness.
Table 3: Reported $p$-values for the sepsis data; values <0.05<0.05 are bolded.
@@ 6.2 Multi-modal Differentiation of Alzheimer’s Disease
Distinguishing cognitive health states using heterogeneous medical data is a central application of multi-modal statistical analysis. In Alzheimer’s disease (AD), differences from cognitively normal (CN) subjects are often subtle and distributed across imaging, metabolic, and clinical biomarkers, making this problem well suited for evaluating methods designed for incomplete multi-source data.
@@
We apply our methods to data from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) (Mueller et al., 2005a). Each participant is identified by a unique Participant Roster ID (RID). Following Balsis et al. (2015), diagnostic groups are defined using Mini-Mental State Examination (MMSE) scores (Folstein et al., 1975): subjects with average MMSE scores in $(24,30]$ are classified as CN, and those with scores $\leq 24$ are classified as AD.
@@
Three modalities are integrated in our analysis: Magnetic Resonance Imaging (MRI), Positron Emission Tomography (PET), and biospecimen data. MRI and PET features are preprocessed by the Center for Imaging of Neurodegenerative Diseases at the University of California, San Francisco, and the Jagust Lab at the University of California, Berkeley, respectively. Biospecimen data, generated by the Proteomics and Metabolomics Shared Resource at Duke University, include serum metabolite measurements such as carnitine and decanoylcarnitine. All data are accessed through the ADNI Analysis Ready Cohort Builder. [^1] https://ida.loni.usc.edu/explore/jsp/search_v2/search.jsp?project=ADNI
After merging modalities by RID, we remove cases with entrywise (random) missingness that do not conform to the modality-level block-wise structure assumed in this work. The resulting dataset consists of $2,555$ subjects with three modalities: MRI, PET, and biospecimen data (BIO), with dimensions $d_{\mathrm{MRI}}=289$, $d_{\mathrm{PET}}=318$, and $d_{\mathrm{BIO}}=518$. Approximately $89.3\%$ of subjects have at least one modality missing, and all seven non-empty modality combinations are observed; see Table 4.
Table 4: Distribution of modality-level missingness patterns in the ADNI data. Here $(S_{\mathrm{MRI}},S_{\mathrm{PET}},S_{\mathrm{BIO}})$ indicates whether each modality is fully observed (11) or missing (0).
$S_{\mathrm{MRI}}$ $S_{\mathrm{PET}}$ $S_{\mathrm{BIO}}$ Count Proportion
0 0 1 592 0.232
0 1 0 375 0.147
0 1 1 642 0.251
1 0 0 119 0.0466
1 0 1 17 0.00665
1 1 0 537 0.210
1 1 1 273 0.107
@@
We evaluate the methods on two datasets: (i) a balanced subset (Sub) with $200$ randomly selected subjects from each group (AD and CN), and (ii) the full dataset (Full) including all $2047$ CN and $507$ AD subjects. For each dataset, inference is conducted using permutation-based ($1000$ permutations) and, when available, asymptotic null distributions.
@@
Table 5 reports the resulting $p$-values. Both BRISE-c and BRISE-v detect highly significant differences between AD and CN subjects across all datasets and inference schemes. In contrast, RISE, MMD, and BD achieve significance only on the full dataset and fail to reject the null in the balanced subset. MMD-Miss yields $p$-values equal to $1$ in all settings, likely due to its inability to exploit cross-pattern information, while MT fails in both datasets. These results are consistent with our simulation findings and demonstrate the practical advantage of BRISE for high-dimensional, block-wise incomplete multi-modal data.
Table 5: Reported $p$-values for the ADNI data; values <0.05<0.05 are bolded.
@@ 7 Discussion and Conclusion
In this work, we propose the general framework BPET for two-sample testing under block-wise missingness without requiring imputation or deletion of observations, accommodating possible MNAR. Further, we incorporate the rank-based RISE test within the proposed BPET framework to account for various types of modalities. We established the finite-sample properties, validity, and consistency of the proposed methods. Their effectiveness was demonstrated through both extensive simulations and real-world applications, and the BRISE exhibited strong empirical performance. Although developed for two-sample testing, BPET can be readily extended to other inferential tasks such as independence testing, clustering, and change-point detection in the presence of block-wise missingness.
As a general framework, we emphasize that BPET is not limited to graph/rank-based methods; a broad class of two-sample test procedures can be applied within the BPET framework. For example, by computing pattern-wise dissimilarities or kernel metrics (Section 2.1), both kernel-based, and distance-based methods can be incorporated into BPET. In multi-modal data, samples from different patterns may have non-overlapping variables or modalities. In the main text, we proposed the zero-filling strategy in Section 2.1; one may also adopt alternative measures, such as rank-based distances (Supplementary M) or embedding all sources into a latent subspace (e.g., via partial multi-view clustering following Li et al. (2014, e.g.); Feng et al. (2024, e.g.)) before computing distances on the aligned projections. An interesting observation from our simulation is that the proposed method leveraging the zero-filling strategy consistently outperforms other dissimilarity measures. This raises a future research question: whether information from samples without shared or overlapping sources/modalities is important, and if so, how to efficiently leverage it. Our work is among the first to explore these important questions.